OpenStudy (anonymous):

*help requested* *fan and medal awarded* Solve for x: 2x - 2 > 4x + 6 x < -4 x > -4 x < 4 x > 4 Simplify the following expression. 5x - 6y + 5z - x - 12y + 2z 6x - 18y + 7z 4x - 18y + 7z 4x - 6y + 7z 6x - 6y + 7z 6 + x + y = x + 6 + y is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

i thinks its a d b

OpenStudy (mathstudent55):

Let's wotk on the first one. 2x - 2 > 4x + 6 You work on a inequality like you work on an equation with one exception. Just as with an equation, you can add the same number to both sides or subtract the same number from both sides. You can also multiply both sides or divide both sides by the same _positive_ number. The difference between an inequality and an equation is that when you multiply or divide both sides of an inequality by a _negative_ number, the inequality sign changes direction.

OpenStudy (mathstudent55):

To start solving this inequality, let's add 2 to both sides 2x - 2 > 4x + 6 +2 +2 -------------- 2x > 4x + 8 Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Now we have 2x > 4x + 8 Now we need to get all terms with x to teh left side, so we'll subtract 4x from both sides 2x > 4x + 8 -4x -4x ----------- -2x > 8 Now we need to divide both sides by -2. here's where you need to be careful with an inequality because when you divide both sides of an inequality by a negative number, the inequality sign changes direction. -2x 8 --- < --- -2 -2 x < -4 That's the answer

OpenStudy (anonymous):

#2 5x - 6y + 5z - x - 12y + 2z first thing u do is add or subtract all the x,y,and z's 5x - 6y + 5z - x - 12y + 2z +x +12 -2 +x +12 -2

OpenStudy (mathstudent55):

Answer is a, which is what you got. Good job!

OpenStudy (anonymous):

okay so for number 2

OpenStudy (mathstudent55):

In the second question, you need to combine like terms. Like terms have the same variable part. 5x - 6y + 5z - x - 12y + 2z 5x - x - 6y - 12y + 5z + 2z 4x - 18y + 7z Here the answer is b

OpenStudy (anonymous):

okay thanks i was a lil confused on that one. okay and for number 3 i pretty much guessed

OpenStudy (mathstudent55):

The last question deals with a property. The commutative property of additon is: a + b = b + a That means you can change the order of an addition. That's what is happening in this question. 6 + x + y = x + 6 + y The only thing that changed from the left side to the right side is 6 + x became x + 6

OpenStudy (anonymous):

oh so i was right ?

OpenStudy (mathstudent55):

Commutative is c. You had d.

OpenStudy (anonymous):

oh sorry i meant to say c. my baaad